OpenStudy (anonymous):

evaluate lim of x approaching infinity of 2xsin(1/x)

OpenStudy (tkhunny):

REALLY BIG hint: \[2x\sin(1/x) = 2\cdot\frac{sin(1/x)}{(1/x)}\]